a first number plus twice a second number is 22. Twice the first number plus the second total 17. Find the numbers.
Answer by oberobic(2304)

x + 2y = 22
2x +y = 17
.
Multiply the 2nd equation by 2 and subtract
.
x + 2y = 22
4x +2y = 34
------------- subtract
-3x = -12
.
x = 4
.
Substitute to find y.
.
4 + 2y = 22
2y = 18
y = 9
.
Check these values using the 2nd equation.
4(4) + 2(9) = 34 ??
16 + 18 = 34
Correct.
.
Answer: The two numbers are 4 and 9.
.
Done.